# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 3.1 Vertical Continuity (Vertical Continual Learning)
Definition. Vertical continuity (or vertical continual learning) has long been studied, either implicitly or explicitly, in existing literature; it involves a sequence of adaptation from general to specific domains and tasks [91, 243, 88, 327, 323]. Along this axis, the training task transitions gradually from general pre-training to downstream tasks, typically undertaken by distinct entities within the production pipeline [231]. Vertical continuity is characterized by a hierarchical structure encompassing data inclusiveness, task scope, and computational resources. Fig. 1 shows a typical pipeline for vertical continuity in LLMs, i.e., “pre-training” $\rightarrow$ “domain-adaptive training” $\rightarrow$ “downstream fine-tuning” [185, 152, 63, 93, 369, 88, 92, 52, 311, 310, 327, 243, 187, 115]:
@@
Throughout the process, the unlabeled domain-specific dataset is smaller in scale than the upstream pre-training phase but larger than the final downstream task fine-tuning phase. This pattern extends to computational resources, team size, and other factors. It is important to note that vertical continuity can involve more than three stages [215, 166, 245, 115]. In real-world applications, during domain-adaptive pre-training, additional layers can be added to accommodate multiple entities, such as various departments with distinct objectives but operating within the same domain.
Figure 2: A diagram showing two different directions of continual learning of LLMs. (a) Vertical Continual Learning of LLMs: in this case, the upstream data distribution usually partially covers the subsequent tasks’ data distribution. (b) Horizontal Continual Learning of LLMs: No constraints on the data distributions are present on horizontal continual learning. The continual LLMs need to handle the challenge of abrupt distributional shifts and longer sequence of training.
@@
Vertical Forgetting. We term the performance degradation on general knowledge of a model undergoing vertical continual learning “vertical forgetting”. As shown in Fig. 2, usually for vertical continual learning, the data distribution of upstream tasks partially covers the downstream, meaning the model might start off at a decent initialization for the subsequent stage of training. However, there are two significant challenges to be addressed to prevent vertical forgetting:
  • Task Heterogeneity. Stemming from the inherent disparity between the formulation of upstream tasks and downstream tasks, task heterogeneity can lead to differences in model structures and training schemes, which has long been recognized as a major hurdle [239, 161, 316, 212, 139]. To mitigate this issue, practitioners often employ methodologies like freezing shared parameters during downstream phases or reformulating downstream tasks to match the structure of pre-training tasks [330, 291, 149, 221, 327, 310].
  • Inaccessible Upstream Data. This challenge arises primarily from varying levels of confidentiality across entities undertaking vertical continual learning. Data collected and curated under different protocols may not be accessible to some downstream entities. This scenario is even more challenging than the strict memory constraint presented in conventional CL (Definition 2.4), as algorithms for latter case rely on access to previous data at specific points for parameter importance measurement [140, 4] or for replay [240, 38, 29, 260]. To address the challenge of inaccessible upstream data, existing methods either use public datasets or generate pseudo-examples to create proxy pre-training dataset [230].
3.2 Horizontal Continuity (Horizontal Continual Learning)
@@
Definition. Horizontal continuity (or horizontal continual learning) refers to continual adaptation across time and domains, a topic extensively explored within the continual learning community. The primary rationale for preserving horizontal continuity lies in the dynamic nature of data distribution over time. To stay updated with these content shifts, an LLM must incrementally learn newly-emerged data. Otherwise, the cost of re-training will become prohibitively expensive and impractical [37, 6, 271, 323]. Empirical evidence has consistently shown that despite their impressive capabilities, LLMs struggle to generalize effectively to future unseen data, particularly in the face of temporal or domain shifts [6, 120, 119, 68]. Additionally, they struggle to retain complete knowledge of past experiences when adapting to new temporal domains, although they do demonstrate a higher level of robustness against catastrophic forgetting [274, 183, 365, 195]. The necessity of employing complex continual learning algorithms to address challenges in LLMs remains an open question. For instance, during large-scale continual pre-training, major institutions can typically afford the storage costs associated with retaining all historical data, rendering memory constraints negligible. Several studies have demonstrated that with full access to historical data, simple sparse replay techniques can effectively mitigate forgetting in large models [277, 274, 255, 228, 81]. In contrast, numerous continual learning studies have showcased superior performance compared to naive solutions, suggesting the importance of continual learning techniques in LLM training [119, 127, 232, 46].
@@
Horizontal Forgetting. We informally define “horizontal forgetting” as the degradation in performance on previous tasks during horizontal continual learning. We informally define “horizontal forgetting” as the performance degradation on the previous tasks when model is undergoing horizontal continual learning. As illustrated in Fig. 2, horizontal continual learning typically involves training stages of similar scales, with potential distributional overlap among their data. In summary, addressing horizontal forgetting presents two main challenges:
@@
  • Longer Task Sequence. Horizontal continual learning ideally involves numerous incremental phases, particularly to accommodate temporal shifts in data distribution. A longer task sequence entails more update steps of the model, leading to inevitable forgetting of previously learned tasks. To address this challenge, researchers employ established continual learning techniques with stronger constraints, such as continual model ensemble [237].
@@
  • Abrupt Distributional Shift. In contrast to vertical continuity, where distributional shifts are often predictable, horizontal continual learning does not impose constraints on the sequential learning tasks’ properties. Evidence suggests that abrupt changes in task distributions can result in significant horizontal forgetting of the model [30, 250].
4 Learning Stages of Continual Large Language Models
@@
Fig. 1 provides an overview of continually learning large language models. Along the axis of vertical continuity, three major layers of modern continual learning emerge. The top layer, Continual Pre-Training (CPT), involves continuous pre-training of LLMs by the supplier on newly-collected data alongside existing data (Section 4.1). As data volume increases, the general capacity of LLMs naturally evolves. The middle layer, Domain-Adaptive Pre-training (DAP), prepares LLMs for domain-specific applications through additional pre-training on domain-specific unlabeled data (Section 4.2). The bottom layer, Continual Fine-Tuning (CFT), targets models for final downstream tasks on the consumer side (Section 4.3). Within continual fine-tuning, we further cover topics including continual instruction tuning (Section 4.3.3), model refinement (Section 4.3.4), model alignment (Section 4.3.5), and multimodal LLMs (Section 4.3.6).
4.1 Continual Pre-Training (CPT)
The recent development of large language models has shattered the glass ceiling in achieving close-to-human levels of natural language understanding and generation. However, effectively adapting these models to the ever-evolving environment remains a fundamental challenge. In Table 1, we outline the basic properties of existing CPT papers.
4.1.1 CPT: Effectiveness and Efficiency
Before delving into the detailed introduction of papers on continual pre-training (CPT), it is important to address two fundamental questions: Firstly, regarding effectiveness, can CPT enhance performance on downstream tasks beyond that of the initial training on a wide range of data domains? Extensive studies, including ELLE [232], DEMix [91], CKL [120], TemporalWiki [119], LLPT [127], and Lifelong-MoE [46], have not only demonstrated the necessity of CPT for improved downstream performance, but also shown that when distributional shifts are gradual [119] or somewhat correlated [91], CPT can enhance model generalization to unseen data.
@@ 4.1.1 CPT: Effectiveness and Efficiency
After confirming the effectiveness of CPT, the second question regarding efficiency arises: given the large number of parameters in the LLM and the size of both old and new data, achieving adaptation and knowledge retention in a computationally efficient manner becomes crucial. Concerning efficiency, most studies focus on techniques for efficient knowledge retention [127, 120, 119, 149], which significantly overlap with the continual learning literature addressing catastrophic forgetting. As mentioned before, these techniques replay [254, 240, 29, 260], parameter regularization [239, 241, 4], and architecture expansion [247, 237, 287]. In contrast to prior approaches that fully utilize emergent data, some studies recognize the impracticality of this approach in real production environments. Instead, they concentrate on further improving the efficiency of adapting to new distributions. For instance, ELLE [232] employs a function-preserved model expansion to facilitate efficient knowledge growth; [6] sub-samples training data based on semantic shift levels to enhance training efficiency; [323] employs a data sampling strategy that encourages novelty and diversity, achieving superior performance to full-data training. Though underexplored, this aspect of efficient adaptation in continual pre-training is poised to become significant, given recent findings emphasizing data quality over quantity for LLM generalization [72, 157, 321, 267].
@@
Table 1: Summary of the existing studies on Horizontal Continual Pre-training of LLMs, where the papers are organized based on their type, where: (i) no continual learning techniques are studied, (ii) continual learning techniques are studied as solely baselines, and (iii) new approaches are proposed, containing some of the continual learning techniques. In the table, Dist. Shift denotes what type(s) of distributional shifts this particular study considers and is dedicated to solve. In the section of Continual Learning Tech., we mainly categorize three types of continual learning techniques that are studied in the paper: rehearsal (Rehearsal), parameter regularization (Param. Reg.), and architecture expansion (Arch. Exp.). We use “✓”, “✗”, and “♣” to denote “deployed in the proposed method”, “not studied in the paper”, and “studied as a baseline method”, respectively; and use “✓” to represent the vocabulary expansion and replacement. It is noteworthy that we do not include naive sequential fine-tuning in this table, as it is universally studied as the important baseline method in all of the papers in the table. The papers with only “♣” [127, 119, 120] means that merely existing CL techniques are studied in them, and the papers with only “✗” [89, 82] means that no CL techniques but special aspects of fine-tuning are studied, e.g., model (re)warming via learning rate scheduling [89].
@@
TimeLMs [175] Temporal 8 RoBERTa
[89] Content 1 Pythia
[82] Language 3 GPT

### v2

@@ 3.1. Vertical Continuity (Vertical Continual Learning)
Definition. Vertical continuity (or vertical continual learning) has long been studied, either implicitly or explicitly, in existing literature. Vertical continuity is characterized by a hierarchical structure encompassing data inclusiveness, task scope, and computational resources. Specifically, the training task transitions gradually from general pre-training to downstream tasks, typically undertaken by distinct entities within the production pipeline (qin2023recyclable, ; gururangan2022demix, ; rongali2021continual, ; guo2023continuous, ; yan2023af, ; xie2023efficient, ). Fig. 1 shows a typical pipeline for vertical continuity in LLMs, i.e., “pre-training” $\rightarrow$ “domain-adaptive training” $\rightarrow$ “downstream fine-tuning” (luo2023biomedgpt, ; li2023cfgpt, ; deng2023learning, ; han2021econet, ; zhou2020pre, ; guo2023continuous, ; gururangan2020dont, ; colombo2024saullm7b, ; wu2023pmc, ; wu2024llama, ; yan2023af, ; rongali2021continual, ; ma2023ecomgptct, ; huang2023lawyer, ):
@@
Figure 1. A high-level overview of the modern pipeline for continually pre-training and fine-tuning LLMs, where two dimensions of continuity are described. Vertical Continuity (or Vertical Continual Learning): LLM training can be vertically divided into three stages: (i) Continual Pre-Training (CPT), (ii) Domain-Adaptive Pre-training (DAP), and (iii) Continual Fine-Tuning (CFT). The main focus is the retention of the LLM’s general knowledge (prevention of vertical forgetting). Horizontal Continuity (or Horizontal Continual Learning): After the LLMs are deployed, the models are continually updated when a new set of data samples becomes available. The primary goal is to prevent horizontal forgetting in a long sequence of tasks.
@@
Throughout the process, the unlabeled domain-specific dataset is smaller in scale than the upstream pre-training phase but larger than the final downstream task fine-tuning phase. This pattern extends to computational resources, team size, and other factors. It is important to note that vertical continuity can involve more than three stages (nijkamp2022codegen, ; lin2023geogalactica, ; rozière2024code, ; huang2023lawyer, ). In real-world applications, during domain-adaptive pre-training, additional layers can be added to accommodate multiple entities, such as various departments with distinct objectives but operating within the same domain.
@@
Vertical Forgetting. We term the performance degradation on general knowledge of a model undergoing vertical continual learning “vertical forgetting”. As shown in Fig. 2, for vertical continual learning, the data distribution of upstream tasks partially covers the downstream, meaning the model might start off at a decent initialization for the subsequent stage of training. There are two significant challenges to be addressed to prevent vertical forgetting:
[…]
Figure 2. A diagram showing two different directions of continual learning of LLMs. (a) Vertical Continual Learning of LLMs: in this case, the upstream data distribution usually partially covers the subsequent tasks’ data distribution. (b) Horizontal Continual Learning of LLMs: No constraints on the data distributions are present on horizontal continual learning. The continual LLMs need to handle the challenge of abrupt distributional shifts and longer sequence of training.
3.2. Horizontal Continuity (Horizontal Continual Learning)
@@
Definition. Horizontal continuity (or horizontal continual learning) refers to continual adaptation across time and domains, a topic extensively explored within the continual learning community. The primary rationale for preserving horizontal continuity lies in the dynamic nature of data distribution over time. To stay updated with these content shifts, an LLM must incrementally learn newly-emerged data. Otherwise, the cost of re-training will become prohibitively expensive and impractical (chaudhry2019efficient, ; amba2021dynamic, ; su2023efficient, ; xie2023efficient, ). Empirical evidence has consistently shown that despite their impressive capabilities, LLMs struggle to generalize effectively to future unseen data, particularly in the face of temporal or domain shifts (amba2021dynamic, ; jang2022towards, ; jang2022temporalwiki, ; dhingra2022time, ). Additionally, they struggle to retain complete knowledge of past experiences when adapting to new temporal domains, although they do demonstrate a higher level of robustness against catastrophic forgetting (tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ). The necessity of employing complex CL algorithms to address challenges in LLMs remains an open question. For instance, during large-scale continual pre-training, large institutions can typically afford the storage costs of retaining all historical data, rendering memory constraints not suitable. Several studies have demonstrated that with full access to historical data, simple sparse replay techniques can effectively mitigate forgetting (thengane2022clip, ; tao2022can, ; scialom2022fine, ; prabhu2023online, ; garg2024tic, ). In contrast, numerous continual learning studies have showcased superior performance compared to naive solutions, suggesting the importance of continual learning techniques in LLM training (jang2022temporalwiki, ; jin2022lifelong, ; qin2022elle, ; chen2023lifelong, ).
@@
Horizontal Forgetting. We informally define “horizontal forgetting” as the performance degradation on the previous tasks when model is undergoing horizontal continual learning. As illustrated in Fig. 2, horizontal continual learning typically involves training stages of similar scales, with potential distributional overlap among their data. In summary, two main challenges need to be addressed for horizontal continual learning of LLMs:
@@
  • Long Task Sequence. Horizontal continual learning ideally involves numerous incremental phases, particularly to accommodate temporal shifts in data distribution. A longer task sequence entails more update steps of the model, leading to inevitable forgetting of previously learned tasks. To address this challenge, researchers employ established continual learning techniques with stronger constraints, such as continual model ensemble (ramesh2021model, ).
@@
  • Abrupt Distributional Shift. In contrast to vertical continuity, where distributional shifts are often predictable, horizontal continual learning does not impose constraints on task properties. Evidence suggests that abrupt changes in task distributions can result in significant horizontal forgetting of the model (caccia2021new, ; sarfraz2023error, ).
4. Learning Stages of Continual Large Language Models
@@
Fig. 1 provides an overview of continually learning LLMs. Along the axis of vertical continuity, three main layers of modern continual learning emerge. The top layer, Continual Pre-Training (CPT), involves continuous pre-training of LLMs by the supplier on newly-collected data alongside existing data (Section 4.1). The middle layer, Domain-Adaptive Pre-training (DAP), prepares LLMs for domain-specific applications through additional pre-training on domain-specific unlabeled data (Section 4.2). The bottom layer, Continual Fine-Tuning (CFT), targets models for final downstream tasks on the consumer side (Section 4.3), where the model needs to be updated after deployment for the specified task.
4.1. Continual Pre-Training (CPT)
4.1.1. CPT: Effectiveness and Efficiency
@@
Before delving into the details of continual pre-training (CPT), it is important to address two fundamental questions: Firstly, regarding effectiveness, can CPT enhance performance on downstream tasks beyond that of the initial training on a wide range of data domains? Extensive studies have not only demonstrated the necessity of CPT for improved downstream performance (qin2022elle, ; gururangan2022demix, ; jang2022towards, ; jang2022temporalwiki, ; jin2022lifelong, ; chen2023lifelong, ), but also shown that when distributional shifts are gradual (jang2022temporalwiki, ; yildiz2024investigating, ) or somewhat correlated (gururangan2022demix, ), CPT can effectively help model generalize to unseen data. The second question is about efficiency: given the large size of LLM’s parameters and data, both old and new, can we achieve adaptation and knowledge retention in a computationally efficient way? Concerning efficiency, most studies focus on techniques for efficient knowledge retention (jin2022lifelong, ; jang2022towards, ; jang2022temporalwiki, ; li2024examining, ), which significantly overlap with the CL literature addressing catastrophic forgetting (schwarz2018progress, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ; rebuffi2017icarl, ; ritter2018online, ; aljundi2018memory, ; rusu2016progressive, ; ramesh2021model, ; wang2022coscl, ). In contrast to prior approaches that fully utilize emergent data, some studies recognize the impracticality of this approach in real production environments. Instead, they concentrate on further improving the efficiency of adaptation. For instance, ELLE (qin2022elle, ) employs a function-preserved model expansion to facilitate efficient knowledge growth; (amba2021dynamic, ) and (xie2023efficient, ) sub-sample training data based on novelty and diversity to enhance training efficiency, achieving superior performance to full-data training. Though currently underexplored, efficient adaptation in continual pre-training is poised to become significant, given recent findings emphasizing data quality over quantity for LLM generalization (du2022glam, ; li2023quality, ; xie2024data, ; soldaini2024dolma, ).
4.1.2. General Observations on CPT
Table 1 summarizes the existing studies on continual pre-training (CPT), and here are some key observations we make about CPT.
4.1.3. Distributional Shifts in CPT
This survey categorizes distributional shifts of CPT into three main types: (i) Language Shift: LLMs sequentially learn different language corpora, e.g., English $\rightarrow$ Chinese (gogoulou2024continual, ; li2024examining, ). (ii) Content Shift: LLMs sequentially learn corpora from different fields, e.g., chemistry $\rightarrow$ biology (gururangan2022demix, ; cossu2022continual, ; jin2022lifelong, ; qin2023recyclable, ; chen2023lifelong, ; gupta2023continual, ). (iii) Temporal Shift: Distributional shifts occur over time, e.g., news in 2021 $\rightarrow$ news in 2022, with a major focus on timestamp-sensitive knowledge retention and update (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ). There is some work continually pre-training LLMs on datasets constructed by errors (zhao2024large, ), re-weighting the samples (chen2024take, ) or tokens (lin2024rho, ), and cannot be properly categorized, we use “Other” to represent them in Table 1.
Language Shift.(gogoulou2024continual, ) focuses on assessing LLMs’ natural ability to learn new languages sequentially (English, Norwegian, and Icelandic). With no explicit CL techniques employed for preventing horizontal forgetting, the study observes consistent positive forward transfer of the knowledge, facilitating new language acquisition regardless of the learning order. Forgetting, on the other hand, emerges as a significant challenge that cannot be mitigated by increasing LLM size. In (li2024examining, ), the degree of forgetting of previously learned language when adapting LLMs to a new language is investigated. Various CL techniques, including parameter freezing, LoRA (hu2021lora, ), and (IA)3 (liu2022few, ), are evaluated across multiple dimensions, including output language, general knowledge retention, and reliability. Preliminary experimental results highlight the non-trivial nature of addressing horizontal forgetting for CPT under the language shift. We argue that research on CPT under language shifts is in its preliminary stages for two main reasons: Firstly, the datasets’ scale, including the number of languages and total token count, remains small. Secondly, specific methods targeting language shifts have yet to be proposed; only basic combinations of existing continual learning techniques have been evaluated.
@@ 4.1.3. Distributional Shifts in CPT
Table 1. Summary of existing studies on Continual Pre-training of LLMs. The papers are organized based on their relation to CL: (i) no CL techniques are studied, (ii) CL techniques are studied as solely baselines, and (iii) new CL approaches are proposed. In the table, Dist. Shift denotes what type(s) of distributional shifts this particular study considers and is dedicated to solve. In the section of Continual Learning Tech., we mainly categorize three types of continual learning techniques that are studied in the paper: rehearsal (Rehearsal), parameter regularization (Param. Reg.), and architecture expansion (Arch. Exp.). We use “✓”, “✗”, and “♣” to denote “deployed in the proposed method”, “not studied in the paper”, and “studied as a baseline method”, respectively. It is noteworthy that we do not include naive sequential fine-tuning in this table, as it is universally studied as the important baseline method in all of the papers in the table. The papers with only “♣” (jin2022lifelong, ; jang2022temporalwiki, ; jang2022towards, ) means that no new but only existing CL techniques are studied in them, and the papers with only “✗” (gupta2023continual, ; gogoulou2024continual, ) means that no CL techniques but special aspects of fine-tuning are studied.
@@
TimeLMs (loureiro2022timelms, ) Temporal 8 RoBERTa
(yildiz2024investigating, ) Content 159 RoBERTa GPT-2
(gupta2023continual, ) Content 1 Pythia
(gogoulou2024continual, ) Language 3 GPT
RHO-1 (lin2024rho, ) Other 1 TinyLlama Mistral